Shift report 20011102 16:00-24:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Fri Nov 02 2001 - 23:58:08 EST

  • Next message: Apache: "Shift report 20011103 00:00-08:00"

    ------------------------------------------------------------
    Date: Fri Nov  2 23:58:08 2001, Shift: 16:00-24:00
    Supervisor: SJS, with: TML
    Log book#10 pages 88 - 89
    
    Shift Summary:
    ------------------------------------------------------------
    1. No beam. Bad beam at the end. Low rate. 
    Started run because they would not inject 
    any new beam for a while. Run 5640 just started.
    
    2. ZDC rates seem to come back when they steer,
    which they tried to during the shift.
    Pedestal runs taken.
    
    3. Had problem with all 3 1440 mainframes. Counld not
    connect to them with brahvo. Turning on and off the 
    terminal server through the webpage fixed the problem.
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Fri Nov 02 2001 - 23:58:39 EST